# ActiveInitSplat: How Active Image Selection Helps Gaussian Splatting

### v1

@@ 4.1 Gaussian process based surrogate model for black-box optimization
Figure 4: The value of (a) $r_{q}$ function, (b) LPIPS, (c) SSIM, and (d) PSNR at each iteration of the selection process for all competing methods in the simulated ‘office 4’ environment [24] where any viewpoint in the 3D space can be selected at each iteration. After 10 iterations, the active selection process of ActiveInitSplat results in higher quality 3D scene representation compared to baselines, along with superior GS rendering performance. In addition, ActiveInitSplat converges faster to the desired rendering performance.
@@
5 Numerical Evaluation
5.1 Implementation details
For demonstration purposes, the 3D representation $\mathcal{P}_{C}(\mathbf{x}_{t};\mathbf{x}_{1},\ldots,\mathbf{x}_{t-1})$ used for the black-box function evaluation $r_{q}(\mathbf{x}_{t};\mathbf{x}_{1}\ldots\mathbf{x}_{t-1})$ upon acquiring $\mathbf{x}_{t}$ at slot $t$, is estimated using COLMAP [22] similarly as in the traditional SfM-based GS model in [12]. However, we note that other alternative techniques for obtaining $\mathcal{P}_{C}(\cdot)$ from 2D images can be readily integrated into our ActiveInitSplat framework; see e.g. a related ablation study in 5.2. For the Gaussian process surrogate model, the kernel functions $\kappa_{s},\kappa_{t}$ used for computing $\kappa$ in (6) are selected to be RBF kernels, which is a typical choice across different GP-based tasks [21]. The kernel hyperparameters are estimated maximizing the marginal likelihood using the sklearn package and are re-fitted at the end of each iteration where a newly acquired $(\mathbf{x}_{t},y_{t})$ pair of camera-view and (noisy) function evaluation becomes available. Note that the kernel function has a significant role in the GP surrogate model and existing works aim to learn the proper kernel form; see e.g. [17, 28, 16, 6], that could be applied in our ActiveInitSplat but exceeds the scope of the current work.
@@ 5.2 Numerical results
To assess the performance of ActiveInitSplat in the 3D scenes described in subsection 5.1, the first step is to corroborate the effectiveness of ActiveInitSplat in identifying camera views from diverse viewpoints that provide high-quality 3D scene representation with increased scene coverage. To that end, Table 1 initially demonstrates the $r_{q}$ function value of all competing methods, that quantifies the quality of a point cloud in terms of density and scene coverage given a specific set of camera views (c.f. section 4). It can be seen in Table 1 that ActiveInitSplat attains a higher $r_{q}$ value compared to the passive selection counterparts in all datasets illustrating the effectiveness of the proposed GP-based black-box optimization technique in actively capturing informative camera views. With the selection of informative 2D images granted, the goal next is to evaluate how this active selection translates into the GS rendering performance for NVS. Table 1 additionally reports the LPIPS, SSIM, and PSNR rendering metrics for all different methods upon completing $T$ iterations of the selection process. It is evident that ActiveInitSplat is the best-performing approach in all datasets and in all rendering metrics implying that the passive methods require a denser selection of images to meet the rendering performance of ActiveInitSplat. Besides quantitative evaluations, Fig. 3 also shows qualitative results of indicative rendered test images compared to the corresponding ground-truth ones in all tested 3D scenes. The annotated dashed yellow boxes in Fig. 3 indicate the differences in the rendered images between all competing methods where it is shown that the rendered images from ActiveInitSplat are closer to the ground-truth ones compared to the passive baselines. Next, we present a number of ablation studies to further highlight the merits of ActiveInitSplat.
@@
A.S.1. How does ActiveInitSplat perform in realistic settings with no constraint on candidate views, i.e. where any viewpoint in the 3D space of interest can be selected?
@@
A.S.2. How does ActiveInitSplat perform when (i) alternative methods (other than COLMAP) are used to generate $\mathcal{P}_{C}(\cdot)$ in the optimization process, and (ii) only a limited number of training images are acquired during the active selection process?
@@ 6 Conclusions
The present work introduces ‘ActiveInitSplat’, a novel approach for actively selecting camera views to capture informative training images, assisting both the initialization and training of GS models. In contrast to existing GS methods that rely on passively pre-selected training images, the proposed framework dynamically acquires 2D images, optimizing density and occupancy-based criteria of the constructed PC from these images. By actively selecting images from diverse viewpoints, ActiveInitSplat ensures comprehensive scene coverage and properly initializes Gaussian functions, conforming them to the ground-truth 3D structure. Numerical tests conducted on both simulated and well-established benchmark datasets highlight the significant advantages of the proposed active selection strategy over conventional passive selection approaches commonly used in existing GS methods.

### v2

@@ 5.1 Implementation details
Figure 4: The value of (a) $r_{q}$ function, (b) LPIPS, (c) SSIM, and (d) PSNR at each iteration of the selection process for all competing methods in the simulated ‘office 4’ environment [27] where any viewpoint in the 3D space can be selected at each iteration. After 10 iterations, ActiveInitSplat results in higher quality 3D scene representation compared to baselines, along with superior GS rendering performance. In addition, ActiveInitSplat converges faster to the desired rendering performance.
@@
Figure 5: Visual comparison of ActiveInitSplat with the passive selection counterparts in the real-world drone-test platform on three indicative test (novel) viewpoints. These test views are captured from diverse angles and include not only the building within the sports field but also the broader surroundings, including background buildings, trees, and roads, which are particularly challenging to capture in such a large-scale environment. It is evident that the rendered images from ActiveInitSplat have substantially improved quality compared to the passive counterpart.
@@ 5.2 Numerical results
To showcase the merits of ActiveInitSplat, we provide numerical results on widely used synthetic and real benchmark datasets, as well as on a real-world drone testing platform, supplemented by additional ablation studies.
5.2.1 Numerical tests on well-known benchmark datasets
@@
To assess the performance of ActiveInitSplat in the 3D scenes described in subsection 5.1, the first step is to corroborate the effectiveness of ActiveInitSplat in identifying camera views from diverse viewpoints that provide high-quality 3D scene representation with increased scene coverage. To that end, Table 1 initially demonstrates the $r_{q}$ function value of all competing methods, that quantifies the quality of a point cloud in terms of density and scene coverage given a specific set of camera views (c.f. section 4). It can be seen in Table 1 that ActiveInitSplat attains a higher $r_{q}$ value compared to the passive selection counterparts in all datasets illustrating the effectiveness of the proposed GP-based black-box optimization technique in actively capturing informative camera views. With the selection of informative 2D images granted, the goal next is to evaluate how this active selection translates into the GS rendering performance for NVS. Table 1 additionally reports the LPIPS, SSIM, and PSNR rendering metrics for all different methods upon completing $T$ iterations of the selection process. It is evident that ActiveInitSplat is the best-performing approach in all datasets and in all rendering metrics implying that the passive methods require a denser selection of images to meet the rendering performance of ActiveInitSplat. Besides quantitative evaluations, Fig. 3 also shows qualitative results of indicative rendered test images compared to the corresponding ground-truth ones in all tested 3D scenes. The annotated dashed yellow boxes in Fig. 3 indicate the differences in the rendered images between all competing methods where it is shown that the rendered images from ActiveInitSplat are closer to the ground-truth ones compared to the passive baselines.
Table 3: LPIPS, SSIM and PSNR values for competing methods in the real-world drone-test platform.
Method LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
3DGS-passive-random 0.4727 0.4566 13.893
ActiveInitSplat 0.3273 0.7024 20.686
Scene ActiveInitSplat (T=12) GS-passive-random (T=12) GS-passive-random (T=20) GS-passive-random (T=30)
Bonsai 5.78 mins 5.22 mins 26.88 mins 164.18 mins
Corn 7.13 mins 6.76 mins 29.31 mins 164.51 mins
Garden 8.48 mins 8.15 mins 62.03 mins 211.36 mins
Table 4: Total runtime comparison between ActiveInitSplat and passive GS alternatives for collecting and processing $T$ camera views for 3D PC generation used for GS initialization.
[…]
CompactGS-passive-random 25.66652 0.3612 0.5669 18.8106
CompactGS-passive-standard 27.83765 0.3164 0.5929 18.7264
ActiveInitSplat & CompactGS 35.43626 0.2649 0.7164 22.5226
bonsai CoRGS-passive-random 25.66652 0.3604 0.5738 18.9091
CoRGS-passive-standard 27.83765 0.3483 0.5808 17.9609
ActiveInitSplat & CoRGS 35.43626 0.2891 0.7046 21.3254
CompactGS-passive-random 0.14839 0.3869 0.5441 17.0439
CompactGS-passive-standard 0.15071 0.3834 0.5513 17.2222
ActiveInitSplat & CompactGS 0.17711 0.3816 0.5902 17.9837
treehill CoRGS-passive-random 0.14839 0.3497 0.5711 18.9142
CoRGS-passive-standard 0.15071 0.3748 0.5542 18.6461
ActiveInitSplat & CoRGS 0.17711 0.3665 0.6023 19.9442
CompactGS-passive-random 0.00314 0.4985 0.7491 18.7709
CompactGS-passive-standard 0.00404 0.4613 0.7691 20.2001
ActiveInitSplat & CompactGS 0.00409 0.4361 0.7934 21.4923
corn CoRGS-passive-random 0.00314 0.4981 0.7505 18.8101
CoRGS-passive-standard 0.00404 0.4722 0.7702 20.2098
ActiveInitSplat & CoRGS 0.00409 0.4402 0.7983 21.6326
Table 5: $r_{q}$, LPIPS, SSIM and PSNR values for all competing methods on real-world 3D scenes using CompactGS and CoRGS architectures and COLMAP for PC generation.
5.2.2 Numerical tests on a real-world drone-test platform
Having demonstrated the benefits of ActiveInitSplat over passive GS counterparts on widely used benchmark datasets with relatively small scenes, in this section we aim to evaluate its rendering performance in a real-world experiment using a drone-test platform over a large scene centered on a building within a sports field. Note that, unlike the benchmark datasets where images are typically captured by humans using cell phones around the scene of interest, capturing comprehensive views in the large-scale drone test platform is impractical when using human cell phones, and constitutes a non-trivial task. In addition to the inherent challenges posed by large-scale environments, 3D rendering on the real-world drone test platform is markedly challenged by: (i) the lower resolution of the extracted images compared to those in the benchmark datasets used in Section 5 of the main paper; (ii) noise in the extracted images introduced by drone movement and adverse weather conditions, such as the strong winds that were encountered while conducting the experiment; and (iii) scene dynamics, including moving vehicles and train, along with tree/plant movements due to wind effects. To that end, this experiment can further highlight the practical utility of active image selection for Gaussian Splatting in more complex, large-scale and challenging real-world environments.
For this experiment, we used a Holybro X500 drone equipped with an SIYI A8 camera to extract images. For drone/camera localization, the reference origin used to obtain the drone position is placed in front of the building. Utilizing this reference origin, the occupancy map/ bounding box required to compute the occupancy score in (4) is constructed with boundaries defined by the outer corners of the sports field. Similarly as in Section 5 of the main paper, for the 3DGS-passive-random baseline, we consider the best-performing in terms of rendering quality out of five independent runs. Note that the ‘3DGS-passive-standard’ baseline is omitted here, as there is no pre-constructed dataset following a specific image acquisition pattern, as in the benchmark datasets. Moreover, adopting a standard circular image collection strategy to form a ‘3DGS-passive-standard’ baseline would certainly require a significantly larger number of images in this large-scale environment, leading to increased computational overhead. For the image collection process used for GS initialization and training, we run both ActiveInitSplat and the passive GS baselines for $T=100$ iterations.
To assess the performance of all competing methods on previously unseen novel (test) views, we illustrate in Fig. 5 the rendered images from ActiveInitSplat and passive GS baseline on three indicative testing viewpoints compared to the ground-truth testing (novel) views. Note that these representative test views are selected from diverse angles and viewpoints, capturing various parts of the scene, including the building within the sports field as well as the broader surroundings such as background buildings, trees, and roads, which are particularly challenging to render accurately. As it can be clearly seen in Fig. 5, the rendered image produced by ActiveInitSplat for the first test viewpoint more closely resembles the ground truth compared to the ‘3DGS-passive-random’ baseline, while in the other two test viewpoints ActiveInitSplat exhibits substantially improved rendering quality performance. Notably, the challening third test view, which predominantly covers regions outside the sports field, is effectively captured by ActiveInitSplat, in contrast to the passive baseline which demonstrates poor rendering performance. This highlights the effectiveness of ActiveInitSplat in selecting informative images with broad scene coverage, facilitating efficient GS initialization and training, and enabling strong generalization across diverse test viewpoints. Lastly, Table 3 reports the LPIPS, SSIM, and PSNR quantitative metrics, showing that ActiveInitSplat achieves substantially superior rendering performance compared to the passive GS baseline. This indicates that the passive approach would require a much denser image collection and processing, and consequently greater computational overhead, to match the performance of ActiveInitSplat.
5.2.3 Additional ablation studies
Next, we pose and aim to answer the following research questions to further highlight the merits of ActiveInitSplat.
@@ 5.2.3 Additional ablation studies
Q1. How does ActiveInitSplat perform in realistic settings with no constraint on candidate views, i.e. where any viewpoint in the 3D space of interest can be selected?
@@
Table 6: Comparison between ActiveInitSplat and active coverage-based 3DGS-FVS baseline on multiple 3D scenes based on the LPIPS, SSIM, and PSNR metrics.
Scene Method LPIPS SSIM PSNR
Flowers ActiveInitSplat 0.4025 0.4191 17.538
3DGS-FVS 0.4317 0.3296 16.455
Bicycle ActiveInitSplat 0.4287 0.4157 17.871
3DGS-FVS 0.4196 0.4239 17.551
Treehill ActiveInitSplat 0.3649 0.5893 18.023
3DGS-FVS 0.3759 0.5569 16.981
Garden ActiveInitSplat 0.1678 0.7536 24.384
3DGS-FVS 0.1745 0.7526 24.162
Bonsai ActiveInitSplat 0.2563 0.7221 22.602
3DGS-FVS 0.2303 0.7613 24.726
Corn ActiveInitSplat 0.4310 0.7920 21.649
3DGS-FVS 0.5155 0.7560 19.450
Q2. How does ActiveInitSplat perform compared to alternative existing active coverage-based image selection strategies?
To further substantiate the advantages of the proposed active image selection strategy in ActiveInitSplat, beyond comparisons with passive baselines, we additionally compare it against an alternative active approach grounded in the farthest view sampling (FVS) principle, which aims to select views that maximize diversity across view space and scene content; see e.g. [34]. It is worth noting that while FVS has been utilized in various settings, including NeRFs, it has not yet been used for GS in the context of 3D scene rendering; yet we integrate FVS into GS as a baseline to underscore the effectiveness of ActiveInitSplat. Table 6 summarizes the rendering performance of the competing approaches, where it can be clearly seen that ActiveInitSplat achieves superior rendering performance on almost all datasets, highlighting its advantages over the active FVS-based alternative.
Q3. How does ActiveInitSplat perform when (i) alternative methods (other than COLMAP) are used to generate $\mathcal{P}_{C}(\cdot)$ in the optimization process, and (ii) only a limited number of training images are acquired during the active selection process?
@@
Runtime comparison. To demonstrate the time efficiency of ActiveInitSplat, Table 4 reports the total runtime for collecting and processing $T$ images for 3D point cloud generation using MASt3R, which is subsequently employed for GS initialization. As shown in Tables 2 and 4, when using $T=12$, ActiveInitSplat attains higher rendering quality than the passive baselines with only a negligible increase in runtime. This implies that more images must be passively collected to meet the rendering performance of ActiveInitSplat. However, increasing the number of images can markedly increase the overall processing cost, as reflected in the runtime results for passive selection with $T=20$ and $T=30$ in Table 4.
Remark 2. It is worth noting that even for the more time-intensive SfM process, its computational overhead can be significantly reduced through incremental updates as new images are acquired online in the active selection process of ActiveInitSplat, instead of re-running the process from scratch.
Q4. How does ActiveInitSplat perform when alternative GS architectures are used?
Here, our aim is to demonstrate the effectiveness of ActiveInitSplat while remaining agnostic to the underlying GS architecture. In addition to the traditional 3DGS model in [12] that relies on the SfM process for initialization, in Q2 we have also tested ActiveInitSplat on a different GS model that alternatively uses the MASt3R pre-trained model in a similar way as in [7]. To further support our claim, we next integrate the proposed ActiveInitSplat into alternative GS architectures, including CoR-GS [38], and Compact-3DGS [15]. Table 5 demonstrates the rendering performance of ActiveInitSplat compared to passive alternatives when using CoR-GS and Compact-3DGS. It is evident that ActiveInitSplat has superior performance compared to the passive counterparts in both CoR-GS and Compact-3DGS, underscoring its versatility and architecture-agnostic advantages across different GS frameworks.
@@ 6 Conclusions
The present work introduces ‘ActiveInitSplat’, a novel approach for actively selecting camera views to capture informative training images, assisting both the initialization and training of GS models. In contrast to existing GS methods that rely on passively pre-selected training images, the proposed framework dynamically acquires 2D images, optimizing density and occupancy-based criteria of the constructed PC from these images. By actively selecting images from diverse viewpoints, ActiveInitSplat ensures comprehensive scene coverage and properly initializes Gaussian functions, conforming them to the ground-truth 3D structure. Numerical tests conducted on both simulated and well-established benchmark datasets highlight the significant advantages of ActiveInitSplat over conventional passive selection approaches commonly used in existing GS methods.
Limitations. A basic limitation of the current work is its adaptivity to dynamic changes in the targeted scenes, since it is designed for static environments. Extending ActiveInitSplat to further account for dynamic 3D scenes belongs to our future research agenda.